OpenStudy (waheguru):

just wanna check my answer please asnwer fas what is the equation of a line in a+b+c form if is passes the points (9,1) (-2,2) PLZZ ANSWER FAST

OpenStudy (waheguru):

plz answer i really need it fast

OpenStudy (anonymous):

\[m=\frac{y2-y1}{x2-x1}\]\[=\frac{1-2}{9+2}\]\[=\frac{-1}{11}\] \[y-y_1=m(x-x_1)\] \[y-1=\frac{-1}{11}x+\frac{9}{11}\] \[y=\frac{-1}{11}x+\frac{20}{11}\]

OpenStudy (waheguru):

u rokk